OpenStudy (anonymous):

sqrt(m^8) simplified is m^4 sqrt(m) right? @radar

OpenStudy (akashdeepdeb):

You can say it is \[\sqrt(m^8)\] \[m^4\] No need of that sqrt(m)!! :) Did you get WHY? :D

OpenStudy (anonymous):

Yes! Thanks!

OpenStudy (unklerhaukus):

\[\sqrt{m^8}=(m^8)^{1/2}=m^{8/2}=m^4\quad\checkmark \]

OpenStudy (akashdeepdeb):

:)